An open cubical container of side `h` is filled with a liquid up to height `h+4`. Mass of liquid inside container is m(neglect atm-pressure) then find a) the maximum constant acc. Of container in horizontal direction so that liquid doesn`t spill out and also force on front wall bottom wall and left wall

Grade:12

1 Answers

Shubham Kumar Gupta Ggps
6 years ago
Conserve area h*h/4=1/2 * h*(h+h`)h`=h/2Tan theta=a/g =h/(h/2) a=2gNow front wall force is mgh/2Back wall force is 2mgBottom is mgh